# Seeding with Differentially Private Network Information

### v2

@@ 5 Simulation Results
Simulations on a synthetic input graph.
To test the performance of our algorithms in a sanitized setup, we generate an input using the Erdős-Rényi random graph model on $200$ nodes with edge probability $0.15$ and run our randomized response and exponential mechanism algorithms with different privacy budgets ($\epsilon=0.1,0.5$, or $1$) and different numbers of influence samples ($m$ ranging from zero to $500$). The independent cascade probability is set to $0.03$. Figure 1 shows that the rate of increase with $m$ slows down with the decreasing privacy budget for both exponential mechanism and randomized response algorithm alike.
Simulations on a real-world dataset.
We use a real dataset representing email interactions between individuals in a large European research institution taken from Stanford Large Network Dataset Collection (Stanford Network Analysis Project 2007, Yin et al. 2017, Leskovec et al. 2007). The nodes represent the individuals and the edges if there is communication between two individuals. This network represents the interaction between $1005$ nodes (individuals) through $25571$ edges. The interactions are mutual, and the graph is undirected. In the simulation setup of Fig. 2, we consider $k=4$ and $k=8$ seeds, the number of influence samples ranges from zero to $1500$, $m\in\{0,100,250,500,750,1000,1250,1500\}$, and the independent cascade probabilities are set to $p=0.0155$ on every edges in ${\mathcal{G}}$ and zero otherwise. The $m=0$ case corresponds to no input information where the choice of $k=4,8$ seeds will be uniformly at random. With no input information at $m=0$, there are no privacy risks to consider.
@@
Figure 2: We test the performance (expected influence $I_{\mathcal{G}}(\cdot)$) of our privacy-preserving seeding algorithms (exponential mechanism and randomized response) on a real-world dataset, at different privacy budgets ($\epsilon$), as number of influence samples ($m$) increases. The figure also shows the non-private, influence sample seeding and the complete-information, deterministic, greedy baselines for comparison. Each point on the plot is average of $50\times 10=500$ trials where we generate $50$ collections of $m$ influence samples, and run our mechanisms on each collection $10$ times. The $50$ realizations capture the randomness in the diffusion process (influence sampling) and for each set of influence samples, we perform $10$ iterations to average out the randomness of the algorithms (randomized response and exponential mechanism). The error bars show the $95\%$ confidence intervals of $500$ trials around the means. Finally, to evaluate the performance of our mechanisms, we use another $2000$ influence samples to estimate the expected size of the spread.
@@
In both Figures 1 and 2, we see that the exponential mechanism outperforms the randomized response mechanism as the latter provides stronger (local) privacy protections. We also observe that low values of privacy budget $\epsilon$ lead to earlier saturation in lower quality expected spread sizes. In those cases, our maximum number of influence samples is not enough to counteract the random noise introduced by the privacy mechanisms, so the value of network information diminishes with the decreasing privacy budge — cf., Eckles et al. (2022, Section 4). Increasing the privacy budget alleviates this effect, as we see a faster increase in the expected spread size with increasing $m$ at higher values of $\epsilon$. In fact, the performance of both mechanisms approaches the original (“non-private”) influence sample greedy seeding algorithm (Algorithm 1), as $\epsilon\to\infty$. For instance, in the randomized response mechanism (Algorithm 4), as the privacy budget increases, the flip probability of the entries in the influence samples $x\in\{0,1\}^{n\times m}$ (Algorithm 4, line 3) decreases, and matrix $C_{\rho,l}$ in Equation 11 is close to the identity matrix, so the performance is similar to the original influence sample greedy seeding algorithm (Algorithm 1). For the exponential mechanism of Section 3.1, at each step the mechanism approximately chooses an optimal seed that has the largest expected marginal contribution to the spread, and a large privacy budget ($\epsilon\to\infty$) sharpens the probability of choosing an optimal seed against others. Figures 1 and 2 also show the deterministic greedy baseline with complete network information to be compared with the exponential mechanism and randomized response algorithms at different $\epsilon$ values with increasing $m$. As we gain better information about the network by increasing the number of influence samples $(m\to\infty)$ and the privacy budget $(\epsilon\to\infty)$, the performance of both algorithms will converge to the deterministic greedy algorithm with complete network information.
@@ 6 Discussions, managerial insights and future directions
Comparison between node-, edge-, and influence-sample DP.
In this paper, we propose a new notion of differential privacy using influence samples adjacency. This notion may be extended to various problems which have underlying network structure. Examples include viral content and rumor cascades on social media, polls in elections from voters who influence each other’s opinions, and medical history of past disease contagions on epidemiological networks.
Note that the input of an influence sample DP algorithm is different from the inputs of node- or edge-DP algorithms – capturing different notions of privacy loss that are not formally comparable; specifically, in influence sample DP an individual may decide whether or not to report their particular value for an influence sample but they cannot change the data generation process. Instead of entry-level adjacency, one may consider individual-level adjacency that two inputs are adjacent if they differ at one row (an individual can decide whether or not to report their entire contagion history). Under this new notion, our analysis for exponential mechanism in Section 3.1 still hold, but we need to strengthen privacy constraints from $\epsilon$ to $\epsilon/m$ where $m$ is the number of samples, because an individual’s change on the influence sample increases from $1$ to $m$. However, as the privacy constraints depend on the number of samples, a large number of samples can no longer ensure that our mechanisms’ performance converges to the original greedy algorithm with full knowledge of the graph. This is inevitable as the new notion prevents our mechanisms from finding high-quality seeds regardless of the number of samples. Indeed, it is impossible to get any meaningful utility for ISDP with individual-level adjacency for similar reasons as edge-DP or node-DP (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg).
@@
In this paper, we see a separation of utility guarantee between node/edge-DP algorithms and influence sample DP algorithms (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg vs LABEL:thm:exp:acc and LABEL:thm:rr:acc). Such separations may be attributed to different notions of adjacency (node, edge, or influence sample adjacency). This observation suggests that differentially private algorithms can have significantly different utility and privacy guarantees by changing the notion of adjacency. However, we are not advocating influence sample DP as superior to node- or edge-DP. On the contrary, one goal of this paper is to bring attention to the role that different notions of adjacency play in discussions of privacy. For instance, because people are connected by social networks, their private data of any forms (census, medical records, etc.) are inevitably affected by the network structure. Those supporting node- and edge-DP may consider that the proper notion of adjacency for two datasets (on census, medical records, etc.) should also depend on the network rather than the traditional Hamming distance. Further discussion can be found in Rezaei and Gao (2019).
@@
Our algorithms have polynomial running times in the size of the graph, with a time complexity of $O(knm)$ and $O(nk^{4}+kn^{2}m)$ for $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ and $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$, respectively, for any constant privacy budget $\epsilon>0$. To improve running times, one may consider improving the iterative sampler for the exponential mechanism in Algorithm 3, or designing a better post-processing step for Line 3 in Algorithm 5. On the other hand, our ISDP randomized response algorithm $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$ is automatically robust against independent randomness noise on influence samples. However, it is unclear how well our ISDP exponential mechanism $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ performs in the presence of noise.
@@ Conclusions.
The correlated nature of social network information has been noted as a vulnerability in DP formulations which provide user-level guarantees with respect to the addition or removal of a single user’s data but ignore information leakage across different users’ records. By focusing on the utility of influence samples for influence maximization, our approach guarantees user’s participation in any given cascade is kept private either locally or centrally and preserves the statistical validity of the cascades for influence estimation and maximization. Subsequently, we provide a rigorous operationalization of action-level privacy by providing users with plausible deniability as to whether or not they have taken part in any particular network cascade (Kifer et al. 2020). This relaxation of user-level DP is ideally suited for the purposes of data collection to inform network intervention design. We expect that formulation and investigation of influence sample DP, with specific attention to the utility of the collected data for network intervention design, should lead to novel generalizations of privacy that are both consequentialist and robust to variations in sampling distributions and randomization noise (Kifer and Machanavajjhala 2014, Rastogi et al. 2009, Liu et al. 2020, Chakrabarti et al. 2022).
[…]
Rahimian and Yu contributed equally and are listed alphabetically. Rahimian is partially supported by the National Science Foundation under Grant No. 2318844. Rahimian also acknowledges support from a Pitt Momentum Funds Grant on Socially Responsible Data Collection and Network Intervention Designs. This research was supported in part by the University of Pittsburgh Center for Research Computing, RRID:SCR_022735, through the resources provided. Specifically, this work used the H2P cluster, which is supported by NSF award number OAC-2117681. Code for reported simulations can be accessed from https://github.com/aminrahimian/dp-inf-max

### v3

@@ 5.1 Simulations on Synthetic Random Graph Models
5.2 Simulations on Empirically-Grounded Sexual Activity Networks
We constructed a dynamic network of men who have sex with men (MSM) using separable temporal exponential random graph models (STERGM) based on data from cross-sectional egocentrically-sampled surveys in the ARTnet study (Anti-Retroviral Therapy Network), conducted in the United States from 2017 to 2019, with 4,909 participants, documenting 16,198 sexual partnerships (Weiss et al. 2020). ARTnet collected data from individuals within the MSM target population, asking them to provide information on the quantity, characteristics, and history of their sexual relationships with other men. In this study, we focus on the San Francisco area due to the availability of data and its representative characteristics within the MSM population (San Francisco has a large and diverse MSM community with a long history of HIV-related public health interventions). Based on this data set, we estimate the STERGM parameters in Appendix 8. Using estimated STERGMs, we simulate sexual behavior among 1,000 individuals over 120 periods of one week each. Each node has fixed attributes, such as gender and age. We grouped every 12 periods as a quarter since our data is simulated on a weekly basis. For each quarter, we randomly selected 300 nodes as initial nodes. From these initial nodes, we identified all individuals connected directly or indirectly through sexual relationships. If node A had a direct sexual relationship with node B and node B subsequently had a relationship with node C, then nodes A and C are considered indirectly connected. This process resulted in 300 influence samples per quarter and a total of 3,000 influence samples. By selecting different initial nodes, we generated another collection of influence samples that are used to evaluate the seed sets that are selected by our algorithm.
Unlike our previous approach, this time we no longer assume the existence of a latent network to construct influence samples through the independent cascade model. Instead, we focus directly on influence samples that better reflect the realities of contact tracing. In reality, it is typically easier to track sexual behavior than to obtain a complete network of new sexual relationships. Using influence samples derived from sexual behavior, our algorithm can identify high-risk nodes for targeted intervention, such as Pre-Exposure Prophylaxis (PrEP). PrEP is a strategy that uses antiretroviral drugs to reduce the risk of HIV infection for people who are not HIV positive but are at risk through sexual activity or injection drug use. Such interventions can be made cost-effective by focusing on individuals at high risk of HIV exposure Kessler et al. (2014), Juusola et al. (2012), Ross et al. (2016). The ARTnet data from 2017 to 2019 coincide with the PrEP uptake period, during which PrEP became increasingly accessible and widely adopted in the MSM community (in 2019, there were approximately 626,000 PrEP users in 77 countries, marking a 69% increase from 2018 (Schaefer et al. 2021)). However, the data required to construct sexual activity or need-sharing networks is highly sensitive, making it crucial to implement proper privacy protections for individuals who share their data for public health interventions.
We ran our privacy-preserving algorithm on influence samples constructed from ARTnet study data. In this case, the chosen seeds are high-risk nodes whose protection provides the most overall benefit in terms of preventing large disease cascades. In the simulation setup of Figure 2, we consider $k=10$ and $k=20$ seeds, the number of influence samples ranges from zero to $3000$, $m\in\{0,500,1000,1500,2000,2500,3000\}$. The case $m=0$ corresponds to the absence of input information where the choice of 10 or 20 seeds will be uniformly random. With no input information at $m=0$, there are no privacy risks to consider.
@@ 5.2 Simulations on Empirically-Grounded Sexual Activity Networks
Figure 2: We evaluate the performance of our privacy-preserving seeding algorithms (exponential mechanism and randomized response) on the MSM dataset by measuring their expected influence ($I_{\mathcal{G}}(\cdot)$) under different privacy budgets ($\epsilon$) as the number of influence samples ($m$) increases. For comparison, the plot also includes non-private baselines from seeding using influence samples and the deterministic greedy with complete network information. Each point on the plot represents the average of 1,000 trials, derived from 50 sets of $m$ influence samples, with each set subjected to 20 iterations of our randomized mechanisms.The error bars display 95% confidence intervals based on these 1,000 trials, but the intervals are too small to be visible. To assess the performance of our mechanisms in each case, we use a separate set of 3000 influence samples to estimate the expected spread size from the seed points that algorithms select.
@@
In both Figures 1 and 2, we see that the exponential mechanism outperforms the randomized response mechanism as the latter provides stronger (local) privacy protections. We also observe that low values of privacy budget $\epsilon$ lead to earlier saturation in lower expected spread sizes. In those cases, our maximum number of influence samples is not enough to counteract the random noise introduced by privacy mechanisms, so the value of network information diminishes with the decreasing privacy budge – cf. Eckles et al. (2022, Section 4). Increasing the privacy budget alleviates this effect, as we see a faster increase in the expected spread size with increasing $m$ at higher values of $\epsilon$. In fact, the performance of both mechanisms approaches the original (“non-private”) influence sample greedy seeding algorithm (Algorithm 1), as $\epsilon\to\infty$. For example, in the randomized response mechanism (Algorithm 4), as the privacy budget increases, the flip probability of the entries in the influence samples $x\in\{0,1\}^{n\times m}$ (Algorithm 4, line 3) decreases and the matrix $C_{\rho,l}$ in Equation 3 approaches the identity matrix, so the performance is similar to greedy non-private seeding on influence samples in Algorithm 1. For the exponential mechanism of Section 3.1, at each step the mechanism chooses the seeds probabilistically with probabilities proportional to the exponential of their marginal contribution to the spread size computed on the influence samples. In this case, increasing the privacy budget ($\epsilon\to\infty$) increases the probability of choosing the seed with the largest marginal contribution. Figures 1 and 2 also show the deterministic baseline for greedy with complete network information to be compared with the exponential mechanism and randomized response algorithms at different $\epsilon$ values with increasing $m$. As we gain better information about the network by increasing the number of influence samples $(m\to\infty)$ and the privacy budget $(\epsilon\to\infty)$, the performance of both algorithms will converge to the deterministic greedy algorithm with complete network information.
@@ 6 Discussions, managerial insights and future directions
Comparison between node, edge, and influence-sample DP.
In this paper, we propose a new notion of differential privacy based on the adjacency of influence samples. This concept can be extended to a wide range of problems that involve an underlying network structure. Examples include viral content and rumor cascades on social networks, polls in elections where voters influence each other’s opinions, and the medical history of past disease contagions in epidemiological networks. It is important to note that the input for an influence sample DP algorithm differs from that of node- or edge-DP algorithms, as they capture different notions of privacy loss that are not formally comparable. Specifically, in the influence sample DP, an individual can decide whether or not to report their particular value for an influence sample, but they cannot alter the data generation process. Sometimes, the underlying network may only exist as an abstract construct resulting from other physically defined relationships, such as the sexual activity network discussed in Section 5. In public health applications such as HIV prevention, where it is difficult to construct a complete sexual activity network due to privacy concerns, contact-tracing data from network cascades may be the only available input that can receive meaningful privacy protection. Our results address entry-level privacy by defining adjacency between cascade entries in the input data set. This can be generalized to individual-level privacy by considering that two inputs are adjacent if they differ in one column, allowing an individual to decide whether or not to report their entire contagion history. By focusing on influence samples rather than the entire network, we ensure that privacy concerns in sensitive contexts, such as HIV prevention, are addressed while still allowing effective data-driven interventions. Under this new notion, our analysis for the exponential mechanism in Section 3.1 still holds, but we need to reduce the privacy budget from $\epsilon$ to $\epsilon/m$, to account for the fact that each individual can modify the entire set of $m$ influence samples. However, as the privacy constraints depend on the number of samples, a large number of samples can no longer ensure that our mechanism’s performance converges to the original greedy algorithm with full knowledge of the graph. This is inevitable as the new notion prevents our mechanisms from finding high-quality seeds regardless of the number of samples. In fact, it is impossible to get any meaningful utility from ISDP with individual-level adjacency for similar reasons as edge or node DP (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg).
@@
In this paper, we see a separation of utility guarantee between node/edge-DP algorithms and influence-sample DP algorithms (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg versus LABEL:thm:exp:acc and LABEL:thm:rr:acc). Such separations may be attributed to different notions of adjacency (node, edge, or influence-sample adjacency). This observation suggests that DP algorithms can have significantly different utility and privacy guarantees by changing the notion of adjacency. Here, we are not advocating the influence sample DP as superior to node or edge DP. In contrast, we would like to bring attention to the role that different notions of adjacency play in discussions of privacy. For instance, because people are connected by social networks, their private data of any form (census, medical records, etc.) are inevitably affected by the network structure. Those supporting node and edge DP may consider that the proper notion of adjacency for two datasets should also depend on the network structure rather than the traditional Hamming distance between the entries. Further discussion can be found in Rezaei and Gao (2019).
@@
Our algorithms have polynomial running times in the size of the graph, with a time complexity of $O(knm)$ and $O(nk^{4}+kn^{2}m)$ for $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ and $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$, respectively, for any constant privacy budget $\epsilon>0$. To improve running times, one may consider improving the iterative sampler for the exponential mechanism in Algorithm 3, or designing a better post-processing step for Line 3 in Algorithm 5. On the other hand, our ISDP randomized response algorithm $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$ is automatically robust against independent random noise in influence samples. In contrast, it is unclear how well our ISDP exponential mechanism $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ performs in the presence of noise.
@@ Conclusions.
The correlated nature of social network information is a vulnerability in DP formulations that provide user-level guarantees with respect to the addition or removal of a single user’s data, but ignore information leakage across different user records. By focusing on the utility of influence samples for influence maximization, our approach guarantees user participation in any given cascade to be kept private either locally or centrally, while preserving the statistical validity of the cascades for influence estimation and maximization. Subsequently, we provide a rigorous operationalization of action-level privacy by providing users with plausible deniability as to whether or not they have participated in any particular network cascade (Kifer et al. 2020). This relaxation of user-level DP is ideally suited for contact tracing purposes to inform public health interventions, such as with the sexual activity network presented in our simulation results. We expect that formulation and investigation of the influence sample DP, with specific attention to the utility of the collected data for the design of network intervention, should lead to novel generalizations of privacy that are consequentialist and robust to variations in sampling distributions and randomization noise (Kifer and Machanavajjhala 2014, Rastogi et al. 2009, Liu et al. 2020, Chakrabarti et al. 2022).
Data Availability
The data used to generate the MSM dynamic sexual behavior dataset originates from the ARTnet study, accessible at https://github.com/EpiModel/ARTnet. Due to confidentiality concerns, this dataset is not publicly accessible. Access to the data requires a Memorandum of Understanding (MOU), which specifies the personnel involved in the analysis and the purposes of the data analysis. The dataset cannot be shared without the consent of the ARTnet Study Principal Investigator (PI), Samuel Jenness at Emory University, as outlined in the MOU. To request access, please contact the PI via email at samuel.m.jenness@emory.edu. Upon review and approval of the MOU, access to the ARTnet dataset will be provided through Github.